OpenStudy (anonymous):

The game of euchre (YOO ker) is played using only the 9s,10s,jack,queen,kings, and aces from a standard desk of cards .Find the probability of being dealt a 5 card hand containing all four suits.

OpenStudy (kinggeorge):

The way I see it, you could solve it like so: Since you need all four suits, you have to choose two cards from one suit, and one from the others. Thus, you have 4 choices for the suit with two cards, two choices of cards from that suit, and six choices from the other suits. Thus, I think the formula should be\[\left(\begin{matrix} 4 \\ 1 \end{matrix} \right)\left(\begin{matrix} 6 \\ 2 \end{matrix} \right)\left(\begin{matrix} 6 \\ 1 \end{matrix} \right)\left(\begin{matrix} 6 \\ 1 \end{matrix} \right)\left(\begin{matrix} 6 \\ 1 \end{matrix} \right) = 4*15*6*6*6=12960\]

OpenStudy (anonymous):

can you explain I still don't get

OpenStudy (anonymous):

total pick 5?

OpenStudy (anonymous):

9s,10s,jack,queen,kings, and aces total 6, how you get 4?

OpenStudy (kinggeorge):

First, you know that only one suit can contain more than one card, and the other three suits must contain only one card. Thus, you have 4 choices for the suit that has 2 cards. (The 4-choose-1 term) Then, from the suit that has 2 cards, you have 6 possible choices, and you choose two of those 6 possible cards. (The 6-choose-2 term) Then, for the three remaining suits, you have to choose only 1 card from each from 6 possible cards. From this you get the 3 different 6-choose-1 terms. The 4's that I use come from the number of suits there are; diamonds, hearts, clubs, spades.

OpenStudy (kinggeorge):

I've chosen 5 cards because one of suits has two cards, so I have 6-choose-2 possibilities for that suit, and then three others have 6-choose-1 possibilities. \(2+1+1+1=5\)

OpenStudy (anonymous):

ty

OpenStudy (kinggeorge):

you're welcome.